### v1

# Spherical Harmonic Oscillators

Abstract.
A linear quantum harmonic oscillator factors into one dimensional oscillators and can be solved using creation and annihilation operators. We consider a spherical analogue (there is an obvious hyperbolic analogue, and natural geometric generalizations). This analogue does not factor, and in fact the two dimensional case is critical and has potential applications to chiral models in 2d quantum field theory.
0. Introduction
A d-dimensional quantum harmonic oscillator is characterized as the unique $SO(d)$ invariant self-adjoint second order differential operator on $\mathbb{R}^{d}$ having principal symbol the Euclidean metric, a Gaussian ground state, and vanishing ground state energy. The oscillator factors as a sum of independent one dimensional oscillators, and it can be solved using annihilation and creation operators; this is true also of q-deformations of the harmonic oscillator (see e.g. [1]).
Consider the stereographic correspondence
$S^{d}\leftrightarrow\mathbb{R}^{d}\cup\{\infty\}:(x_{0},x_{1},...,x_{d})\leftrightarrow\frac{(x_{1},...,x_{d})}{1-x_{0}}$
The round metric in stereographic coordinates is given by the formula
$(ds)_{S^{n}}=2\frac{\sqrt{\sum_{j=1}^{d}(dx_{j})^{2}}}{1+|x|^{2}}$
the volume form is given by
$dV_{S^{d}}=2^{d}(1+|x|^{2})^{-d}dV_{\mathbb{R}^{d}}=2^{d}\frac{r^{d-1}}{(1+r^{2})^{d}}drdV_{S^{d-1}}$
and the nonnegative Laplace operator is given by
$\Delta_{(S^{d})}=-\frac{1}{4}(1+r^{2})^{2}\sum_{k=1}^{d}(\frac{\partial}{\partial x_{i}})^{2}$
where $r=|x|$.
Given $\omega\geq 0$, the spherical harmonic oscillator is the differential operator
$L_{\omega}:=\Delta_{S^{d}}+\omega^{2}r^{2}$
This is essentially self-adjoint on compactly supported smooth functions on $\mathbb{R}^{d}$ (see Section 1). It is routine to find the spectrum of this operator, it does not factor, and one cannot use annihilation and creation operators. This begs the question of why we bother to write a paper about it.
A Gaussian is the ground state of a linear harmonic oscillator. Consider (the Cauchy type density) $(1+r^{2})^{-\omega}$. Is this the ground state of a natural geometric operator? If $d=2$, then the answer is yes, there is a unique $SO(d)$ invariant self-adjoint second order differential operator having principal symbol determined by the rotationally invariant metric for the d-sphere and having the ground state $(1+r^{2})^{-\omega}$, with vanishing ground state energy:
$L_{\omega}-\omega:=\Delta_{S^{d}}+\omega^{2}r^{2}-\omega$
This Cauchy type ground state arises in a natural way (when $d=2$) in connection with chiral models in 2d quantum field theory (see Section 5 of [7]).
The story is a little more complicated, in an interesting way, when we consider the dimension $d$ as a parameter. In general the quadratic potential is quite singular from the point of view of the original $d$-sphere, and breaks the $O(d)$ symmetry down to $O(d-1)$. The domain of the operator depends sensitively on $\omega$, e.g. $1$ is in the domain when $\omega=0$, and not when $\omega>0$. Various quantities depend discontinuously on the dimension parameter $d$, and $d=2$ is critical. This is essentially because the principal part of the form corresponding to the operator is conformally invariant when $d=2$. By considering general $d$ we learn something about the $d=2$ case.
There are variations on this theme, e.g. with the hyperbolic disk in place of the sphere. It is not currently clear if we learn anything new from these variations, or if we are even asking the right questions. Consequently we will put this off to another time.
0.1. Plan of the Paper
In Section 1 we use standard arguments to show that $L_{\omega}$ (with appropriate domain) is a nonnegative self-adjoint operator with compact resolvent (hence discrete spectrum).
In Section 2 we find the groundstate and observe how various formulas depend discontinuously on the dimension parameter; $d=2$ is critical.
In Sections 3 and 4 we find the spectrum. We have separated out the $d=2$ because of its importance and relative simplicity.
1. Self-Adjointness
Recall the following three examples from page 331 of [8] of unbounded operators $(T_{k},D(T_{k}))$ in the real Hilbert space $H:=L^{2}([0,1],dx;\mathbb{R})$: $D(T_{1})$ consists of absolutely continuous functions $f$ on $[0,1]$ such that $f^{\prime}\in L^{2}$,
$D(T_{2}):=\{f\in D(T_{1}):f(0)=f(1)\},\text{ and }D(T_{3}):=\{f\in D(T_{1}):f(0)=f(1)=0\}$
and in each case $T_{k}f=\frac{1}{i}f^{\prime}$. Then
$T_{1}^{*}=T_{3},\qquad T_{2}^{*}=T_{2},\qquad T_{3}^{*}=T_{1}$
Heuristically the spectrum of $T_{1}$ consists of all $\lambda\in\mathbb{C}$ with corresponding eigenfunction $f(x):=e^{i\lambda x}$, the spectrum of the self-adjoint operator $T_{2}$ is $2\pi\mathbb{Z}$, and the spectrum of $T_{3}$ is empty. It is important that our spherical oscillators are self-adjoint, to avoid the ”pathology” that occurs in these examples.
An efficient way to define a non-negative self-adjoint operator is to use symmetric forms, see page 11 of [3]. The Hilbert space is $H:=L^{2}(dV_{S^{d}},\mathbb{R})$. The form is
$Q_{\omega}(f,g):=\int(df\wedge*dg+\eta^{2}r^{2}fgdV_{S^{d}})$
where * denotes the star operator for $S^{d}$ (which is conformally invariant precisely when $d=2$), and $Q=Q_{\omega}$ is defined for $f,g$ in the dense subspace $E\subset H$ for which $(Qf,f)$ and $Q(g,g)$ are finite.
Definition 1.
(following [3])
$D(L_{\omega}):=\{g\in E:f\to Q(f,g)\text{ is continuous on }H\}$
and $Q(f,g)=\langle f,L_{\omega}g\rangle_{H}$.
Theorem 1.1.
$L_{\omega}$ is self-adjoint and the resolvent is compact (more precisely in the Dixmier Schatten class $\mathcal{L}_{d-1}^{+}$).
The first statement follows from the Form Representation Theorem 2.1, page 12 of [3]. Note that $L_{\omega}$ as a differential operator is elliptic in $\mathbb{R}^{d}$ and equal to the self-adjoint $L_{\omega}$ on sufficiently differentiable functions. The second statement follows from the fact that the inverses are pseudo-differential operators of order -2. Note that in the case $d=2$ the inverses are actually in the Dixmier Schatten class $\mathcal{L}_{1}^{+}$.
2. The Ground State
The following shows how the ground state energy depends discontinuously on $\omega$ and the dimension parameter.
Theorem 2.1.
If $\omega>0$, then for any dimension $d$, the ground state for $L_{\omega}$ is
$(1+r^{2})^{-\frac{d-2+((d-2)^{2}+16\omega^{2})^{1/2}}{4}}$
and the ground state energy is
$\lambda^{(d)}=d\frac{d-2+((d-2)^{2}+16\omega^{2})^{1/2}}{8}$
where we choose the positive square root in both formulas.
If $\omega=0$, then the above formulas are valid, provided that for $d=1,2$, we choose the positive square roots, and for $d>2$, we choose the negative square roots.
Proof.
Define $\delta(r):=(1+r^{2})^{-\eta}$, $\eta>0$. Then
$L_{\omega}\delta=\frac{(-2\eta^{2}+(d-2)\eta+2\omega^{2})r^{2}+d\eta}{2}\delta$
For $\delta$ to be an eigenfunction (with eigenvalue $d\eta/2$) it is necessary that $-2\eta^{2}+(d-2)\eta+2\omega^{2}=0$, i.e.
$\eta=\frac{d-2\pm((d-2)^{2}+16\omega^{2})^{1/2}}{4}$
This must be nonnegative, vanish when $\omega=0$, and $\delta$ must also be in the domain, in particular
[…]
These conditions determine the signs of the square roots that appear in the statement of the theorem. ∎
Theorem 2.2.
Suppose $\omega>0$. There is a unique $SO(d)$ invariant self-adjoint second order differential operator having principal symbol determined by the rotationally invariant metric for the 2-sphere and having the ground state $(1+r^{2})^{-\omega}$, with vanishing ground state energy,
$L:=-\frac{1}{4}(1+r^{2})^{2}\left((\frac{\partial}{\partial r})^{2}+\frac{d-1}{r}\frac{\partial}{\partial r}+\frac{1}{r^{2}}\Delta_{(S^{d-1})}\right)+\left(\omega^{2}+(2-d)\omega\right)r^{2}-\frac{1}{2}d\omega$
Proof.
We are led to this formula by the calculation in the preceding proof (with $\eta=\omega$).
Imagine adding lower order terms to $L_{\omega}$. The corresponding quadratic form, in terms of the radial parameter $r$ (we can ignore $\theta$ because of rotational invariance), would (because of self-adjointness) be of the form
[…]
$=\int_{r=0}^{\infty}(b(r)f^{\prime}(r)g(r)\frac{r}{(1+r^{2})^{2}}-\frac{\partial}{\partial r}(b(r)f(r)\frac{r}{(1+r^{2})^{2}})dr$
This calculation shows that for the differential operator, the derivative terms cancel, and this only adds zeroth order terms. The fact that we know the lowest energy state implies that $L$ is determined up to a constant. The fact the ground state energy vanishes determines the constant. ∎
3. Spectrum: The Two Dimensional Case
Theorem 3.1.
Suppose $\omega\geq 0$.
(a) The ground state for $L_{\omega}$ is
$(1+r^{2})^{-\omega}$
with corresponding eigenvalue $\omega$.
(b) The spectrum is $\omega$, with multiplicity one, and
$\lambda_{m,n}=m^{2}+m+mn+\frac{1}{2}(n^{2}+n)-\frac{1}{2}(2m+n+1)\sqrt{4\omega^{2}+n^{2}}$
$m,n\geq 0$, $(m,n)\neq 0$, with multiplicity 2 (We are not able to rule out the possibility that these values have redundancy, which would increase the multiplicity, see (c))
(c) If $\omega=0$, then the spectrum is $N(N+1)$ with multiplicity $12N+$ 1, $N=0,1,...$.
Proof.
We have already proven (a) (We will add an abstract geometric argument in Remark Remark following the proof, to explain how we were led to the operator from the groundstate).
To find eigenvalues and eigenfunctions, we use separation of variables. Suppose we consider an eigenfunction of the form $g(r)f(\theta)$. Then
$-\frac{1}{4}(1+r^{2})^{2}\left(g^{\prime\prime}(r)f(\theta)+\frac{1}{r}g^{\prime}(r)f(\theta)+\frac{1}{r^{2}}g(r)f^{\prime\prime}(\theta)\right)+\omega^{2}r^{2}g(r)f(\theta)=\lambda g(r)f(\theta)$
Then
$-(1+r^{2})^{2}\left(\frac{g^{\prime\prime}(r)}{g(r)}+\frac{1}{r}\frac{g^{\prime}(r)}{g(r)}+\frac{1}{r^{2}}\frac{f^{\prime\prime}(\theta)}{f(\theta)}\right)+4\omega^{2}r^{2}=4\lambda$
and
$\frac{r^{2}}{(1+r^{2})^{2}}\left(-(1+r^{2})^{2}(\frac{g^{\prime\prime}(r)}{g(r)}+\frac{1}{r}\frac{g^{\prime}(r)}{g(r)})+4\omega^{2}r^{2}-4\lambda\right)=\frac{f^{\prime\prime}(\theta)}{f(\theta)}$
This implies that $f(\theta)$ is a combination of $cos(n\theta)$ and $sin(n\theta)$, where $n$ is a nonnegative integer, and
$r^{2}g^{\prime\prime}(r)+rg^{\prime}(r)+\left(-n^{2}+4(1+r^{2})^{-2}(-\omega^{2}r^{4}+\lambda r^{2})\right)g(r)=0$
This has regular singular points and is equivalent to the hypergeometric equation. The indicial equation is $\alpha^{2}-n^{2}=0$ and $g$ has the form
$g(r)=r^{\pm n}\sum_{l=0}^{\infty}g_{l}r^{l}$
The relevant solution involves $+n$, so that $g$ is regular at $r=0$. One finds
$g(r)=\frac{r^{n}}{(1+r^{2})^{(\sqrt{4\omega^{2}+4\lambda+1}-1)/2}}F(a,b,n+1,-r^{2})$
where $F$ denotes the hypergeometric function with
$a=\frac{1}{2}(-\sqrt{4\omega^{2}+4\lambda+1}+n+1-\sqrt{4\omega^{2}+n^{2}})$
and
$b=\frac{1}{2}(-\sqrt{4\omega^{2}+4\lambda+1}+n+1+\sqrt{4\omega^{2}+n^{2}})=a+\sqrt{4\omega^{2}+n^{2}}$
A necessary condition for $\lambda$ to be an eigenvalue is that $g(r)$ is square integrable on $(0,\infty)$ with respect to $(1+r^{2})^{-2}rdr$. Thus $a$ or $b$ must be a non-positive integer. As above $a<b$.
Examples.
Suppose $\omega=0$. In this case $b=a+n$. Suppose $b=-m$, a non-positive integer. In this case $\lambda=\lambda_{m,n}=(m+n)(m+n+1)$. Square integrability holds. Now consider the possibility that $a=-m$, a non-positive integer and $b=a+n=-m+n>0$, i.e. $n>m$. In this case $\lambda=m(m+1)$. For $g(r)$ to be square integrable it is necessary that
2($12(\sqrt{1+4\lambda}-1)-2n-4m+3>$ 1
or
$1+4\lambda=1+4m(m+1)>(n+2m)^{2}=n^{2}+4mn+4m^{2}$
or
$1+4m(1)>(n+2m)^{2}=n^{2}+4mn$
Square integrability fails. So we need both a and b to be non-positive. If we fix $\lambda=N(N+1)$, then the multiplicity is $1+2N$. For if $\lambda=\lambda_{m,n}$, when $n=0$ we can only choose $cos(n\theta)=1$ and when $n>0$ we can choose $cos(n\theta)$ or $sin(n\theta)$.
Now suppose $\omega>0$.
Suppose that $b=-m$, a non-positive integer. Since $a<b$, $F(a,b,c,-r^{2})$ is a polynomial of degree $m$ in $z=-r^{2}$. In this case
$\lambda=\lambda_{m,n}=m^{2}+m+mn+\frac{1}{2}(n^{2}+n)+\frac{1}{2}(2m+n+1)\sqrt{4\omega^{2}+n^{2}}$
We first find the condition for $g\in L^{2}((1+r^{2})^{-1}rdr)$. Since $F(a,b,n+1,-r^{2})$ is a polynomial of degree $2m$ in $r$, $g\in L^{2}$ if and only if
$2(\sqrt{4\omega^{2}+4\lambda+1}-1)-2n-4m)+3>1$
or
$4\omega^{2}+4(m^{2}+m+mn+\frac{1}{2}(n^{2}+n)+\frac{1}{2}(2m+n+1)\sqrt{n^{2}+4\omega^{2}})+1>(n+2m)^{2}$
This condition is always satisfied, so $g(r)f(\theta)$ is in the Hilbert space. To see that it is actually in $D(L_{\omega})$, for $F\in C^{\infty}(\mathbb{R}^{2})$ having compact support, we use integration by parts. This is justified, because the functions are smooth and $F$ vanishes outside of a compact set. Consequently
$\int(dF\wedge*d(gf)+\omega^{2}r^{2}Fgf)=\int F((\Delta+\omega^{w}r^{2})(gf))dV_{S^{d}}=\langle F,\lambda gf\rangle_{H}$
This shows that $H\to\mathbb{R}:F\to Q(F,gf)$ is continuous, hence $gf\in D(L_{\omega})$. Thus $\lambda_{m,n}$ is actually an eigenvalue.
Suppose that $a=-m$ and $b=a+\sqrt{4\omega^{2}+n^{2}}>0$. In this case
$\lambda=\lambda_{m,n}^{\prime}=m^{2}+m+mn+\frac{1}{2}(n^{2}+n)-\frac{1}{2}(2m+n+1)\sqrt{n^{2}+4\omega^{2}}$
Again $F(a,b,n+1,-r^{2})$ is a polynomial of degree $2m$ in $r$. Thus $g\in L^{2}$ if and only if
$12(\sqrt{4\omega^{2}+4\lambda+1}-1)-2n-4m+3>$ 1
or
$02\sqrt{4\omega^{2}+4\lambda+1}-2n-4m>$ 0
or
$4\omega^{2}+4\lambda+1>(n+2m)^{2}$
or
$4\omega^{2}+4(m^{2}+m+mn+\frac{n^{2}+n}{2}-\frac{1}{2}(2m+n+1)\sqrt{n^{2}+4\omega^{2}})+1>(n+2m)^{2}$
or
$4\omega^{2}+4m+(n+1)^{2}>2(2m+n+1)\sqrt{n^{2}+4\omega^{2}}$
or
$(1-n^{2}-4\omega^{2})(16m^{2}+16mn+3n^{2}-4\omega^{2}+8m+4n+1)>0$
Suppose $1-n^{2}-4\omega^{2})\geq 0$. Since we are assuming $\omega>0$, this implies $n=0$. But then $b=-m+2\omega\leq 0$, which contradicts our assumption that $b>0$. Thus $1-n^{2}-4\omega^{2})<0$.
By way of contradiction, suppose that $\lambda_{m,n}^{\prime}$ is an eigenvalue. Then
$16m^{2}+16mn+3n^{2}+8m+4n+1<4\omega^{2}$
or
$(4m+1)^{2}+(3n+4)n<4\omega^{2}$
We also need the eigenvalue to be nonnegative, i.e.
$m^{2}+m+mn+\frac{1}{2}(n^{2}+n)-(2m+1+n)\sqrt{n^{2}+4\omega^{2}}\geq 0$
We therefore obtain two conditions
$n^{2}+4\omega^{2}>(4m+1)^{2}+(2n+4)n$
and
$\frac{m^{2}+m+mn+\frac{1}{2}(n^{2}+n)}{2m+1+n}\geq\sqrt{n^{2}+4\omega^{2}}$
These two conditions do not have any solutions in $(m,n)$. Thus $\lambda_{m,n}^{\prime}$ is not an eigenvalue.
This completes the proof.
Remark.
When $d=2$, the ground state calculation, and the origin of the operator $L_{\omega}$, can be cast in more abstract terms. Suppose that $L$ is a holomorphic Hermitian line bundle with canonical connection $\nabla$, where $\nabla^{(0,1)}=\overline{\partial}$. If $s$ is a holomorphic section, then
$\partial(s,s)=\theta(s,s)$
and
¯$\overline{\partial}\partial(s,s)=\overline{\partial}(\theta)(s,s)-\theta\wedge\overline{\theta}(s,s)$
(see page 73 of [5]). Thus
¯$\overline{\partial}\partial(s,s)+\theta\wedge\overline{\theta}(s,s)=\Theta(s,s)$
where $\Theta$ is the curvature, a (1,1) form. In our case the line bundle is the kth power of the dual of the canonical bundle (or the 2kth power of the generating positive line bundle on $\mathbb{P}^{1}$, the dual of the tautological bundle). The section $s$ is a canonical (highest weight) holomorphic section such that $(s,s)=(1+r^{2})^{-2k}$ and the curvature is a multiple (essentially the ground state energy) of the area form on the base (because of homogeneity).
This point of view leads to one possible generalization, involving Kahler-Einstein manifolds (where the Ricci form is a multiple of the symplectic form on the base).
4. Spectrum: The General Case
If $\omega=0$, then it is well-known that the spectrum is $N(N+d-1)$ with multiplicity
$\left(\begin{matrix}(N+d\\ N\end{matrix}\right)-\left(\begin{matrix}(N+d-2\\ N-2\end{matrix}\right),\qquad N=0,1,...$
. In particular the ground state energy is zero.
Theorem 4.1.
Suppose that $\omega>0$.
(a) The ground state is
$\delta(r):=(1+r^{2})^{-\frac{d-2+((d-2)^{2}+16\omega^{2})^{1/2}}{4}}$
with ground state energy
$\lambda^{(d)}=d\frac{d-2+((d-2)^{2}+16\omega^{2})^{1/2}}{8}$
where we choose the positive square root in both formulas.
(b) The spectrum is $\omega$, with multiplicity one, and
$\lambda_{m,n}=m^{2}+m+mn+\frac{1}{2}(n^{2}+n)-\frac{1}{2}(2m+n+1)\sqrt{4\omega^{2}+n^{2}}$
$m,n\geq 0$, $(m,n)\neq 0$, with multiplicity
$\left(\begin{matrix}(N+d-1\\ N\end{matrix}\right)-\left(\begin{matrix}(N+d-3\\ N-3\end{matrix}\right)$
where $N=n(n+d-2)$ (As in the $d=2$ case, we are not able to rule out the possibility that these values have redundancy, which would increase the multiplicity, as in the case $\omega=0$)
Proof.
We proved (a) in Section 2 and observed this depends discontinuously on $\omega$ and the dimension. This will be important in Remark Remark below.
(b) To find eigenvalues and eigenfunctions, we use separation of variables. Suppose we consider an eigenfunction of the form $g(r)f(\theta)$, where (abusing notation) we are writing $f(\theta)$ for a function that is defined on the $d-1$ sphere. Then
$-\frac{1}{4}(1+r^{2})^{2}\left(g^{\prime\prime}(r)f(\theta)+\frac{d-1}{r}g^{\prime}(r)f(\theta)+\frac{1}{r^{2}}g(r)\Delta_{S^{d-1}}f(\theta)\right)+\omega^{2}r^{2}g(r)f(\theta)=\lambda g(r)f(\theta)$
Thus
$-(1+r^{2})^{2}\left(\frac{g^{\prime\prime}(r)}{g(r)}+\frac{d-1}{r}\frac{g^{\prime}(r)}{g(r)}+\frac{1}{r^{2}}\frac{\Delta_{S^{d-1}}f(\theta)}{f(\theta)}\right)+4\omega^{2}r^{2}=4\lambda$
and
$\frac{r^{2}}{(1+r^{2})^{2}}\left(-(1+r^{2})^{2}(\frac{g^{\prime\prime}(r)}{g(r)}+\frac{d-1}{r}\frac{g^{\prime}(r)}{g(r)})+4\omega^{2}r^{2}-4\lambda\right)=\frac{\Delta_{S^{d-1}}f(\theta)}{f(\theta)}$
This implies that $f(\theta)$ is an eigenfunction for $\Delta_{S^{d-1}}$. The corresponding eigenvalue is of the form $n(n+d-2)$, where $n$ is a nonnegative integer, and
$r^{2}g^{\prime\prime}(r)+(d-1)rg^{\prime}(r)+\left(-n(n+d-2)+4(1+r^{2})^{-2}(-\omega^{2}r^{4}+\lambda r^{2})\right)g(r)=0$
This has regular singular points and is equivalent to the hypergeometric equation. The indicial equation is
$\alpha^{2}+(d-2)\alpha-n(n+d-2)=0$
for which the roots are $\alpha=n$ and $\alpha=-(n+d-2)$. For $g(r)$ to be regular at $r=0$, $g(r)$ must be of the form
$g(r)=r^{n}\sum_{l=0}^{\infty}g_{l}r^{l}$
We find
$g(r)=r^{n}\left(1+r^{2}\right)^{\frac{1}{2}-\frac{\sqrt{4\omega^{2}+4\lambda+1}}{2}}F(a,b,n+\frac{d}{2},-r^{2})$
where $F$ denotes the hypergeometric function, and
$a=-\frac{\sqrt{4\omega^{2}+4\lambda+1}}{2}+\frac{n}{2}+\frac{d}{4}-\frac{\sqrt{(2n+d-2)^{2}+16\omega^{2}}}{4}$
$b=-\frac{\sqrt{4\omega^{2}+4\lambda+1}}{2}+\frac{n}{2}+\frac{d}{4}+\frac{\sqrt{(2n+d-2)^{2}+16\omega^{2}}}{4}$
For $\lambda$ to be an eigenvalue, $g(r)$ must be square integrable on $(0,\infty)$ with respect to $(1+r^{2})^{-d}r^{d-1}dr$. A necessary condition is that $a$ or $b$ is a non-positive integer.
Suppose that $a<b=-m$, a non-positive integer. In this case $F(a,b,n+\frac{d}{2},-r^{2})$ is a polynomial of degree $m$ in $z=-r^{2}$. In this case
$\lambda=\lambda_{m,n}=m^{2}+\frac{md}{2}+mn+\frac{n^{2}+n(d-1)-2(d-2)}{2}+\frac{(4m+2n+d)\sqrt{(2n+d-2)^{2}+16\omega^{2}}}{8}$
Remark.
(a) This agrees with our formula for $d=2$.
(b) We are assuming $\omega>0$. If we insert $\omega=0$, this formula simplifies to
$\lambda_{m,n}=(m+n+\frac{d}{2})(m+n+\frac{d-2}{2})$
This is not the correct form for the eigenvalues when $d\neq 2$. The problem is that we have to account for the discontinuity in the ground state energy. Suppose $\omega>0$. If we normalize the ground state energy to be zero, then we are considering
$\lambda_{m,n}-d\frac{d-2+((d-2)^{2}+16\omega^{2})^{1/2}}{8}$
where we take the positive square root. When $\omega\downarrow 0$, this simplifies to $(m+n)(m+n+d-1)$, which is the correct form of the $\omega=0$ eigenvalues.
For $\lambda_{m,n}$ to be an eigenvalue it is necessary that $g\in L^{2}((1+r^{2})^{-d}r^{d-1}dr)$. Now $F(a,b,n+\frac{d}{2},-r^{2})$ is a polynomial of degree $2m$ in $r$. Thus $g\in L^{2}$ if and only if
$12(\sqrt{4\omega^{2}+4\lambda+1}-1)-2n-4m+2d-(d-1)>$ 1
or
$02(\sqrt{4\omega^{2}+4\lambda+1}-1)-2n-4m+d>$ 0
or
$4\omega^{2}+4\lambda+1>(n+2m-\frac{d-2}{2})^{2}$
or
$14\omega^{2}+4(m^{2}+\frac{md}{2}+mn+\frac{n^{2}+n(d-1)-2(d-2)}{2}+\frac{(4m+2n+d)\sqrt{(2n+d-2)^{2}+16\omega^{2}}}{8})+$ 1
$>(n+2m-\frac{d-2}{2})^{2}$
This condition is always satisfied, so $g(r)f(\theta)$ is in the Hilbert space. We use the same integration by parts argument to see that $g(r)f(\theta)\in D(L_{\omega})$, which implies $\lambda_{m,n}$ is an eigenvalue.
As in the $d=2$ case we rule out $\lambda$ as an eigenvalue when $a=-m$ and $b>0$. ∎
4.1. Partition and Zeta Functions
In the case $d=2$ it is well=known that the partition function $tr(e^{-tL_{\omega=0}})$ can be expressed in terms of the theta function with characteristics, $\theta[\begin{matrix}1\\ 1\end{matrix}](z,\tau)$ (or $\theta_{1}$), where $\theta^{\prime}$ denotes differentiation with respect to $z$ (see chapter 2, section 8 of [4]): If $\omega=0$ and $q=e^{-t}=e^{i\pi\tau}$, $\tau=t/i\pi$, then
$tr(e^{-tL_{\omega=0}})\sum_{N=0}^{\infty}(2N+1)q^{N^{2}+N}=\theta^{\prime}[\begin{matrix}1\\ 1\end{matrix}](\pi,8\tau)$
(and this can be expressed as an infinite product using the Jacobi product formula). I do not know if there are similar expressions when $d\neq 2$.
In the case $\omega=0$, Vardi ([9] with some corrections as in [6] and [2]) showed how to express the zeta function for $L_{\omega=0}$ in terms of the Riemann zeta function, and he used this to evaluate the zeta determinant.
It is trivial that (for $d=2$, and for other $d$ as well)
$tr(e^{-tL_{\omega}})=1+2\sum_{m,n\geq 0}e^{-t(m^{2}+m+mn+\frac{1}{2}(n^{2}+n)-\frac{1}{2}(2m+n+1)\sqrt{n^{2}+4\omega^{2}}}\leq tr(e^{-tL_{0}})$
so the partition function is convergent for any $\omega$. It would seem unlikely that this can be expressed in terms of standard functions, even for special values of $\omega$.
Since the potential is singular, it does not seem to follows from standard arguments that there is a zeta function with an meromorphic continuation. However it should be possible to given an ad hoc argument, by finding an asymptotic expansion for the partition function.