### v2

# On a Spherical Analogue of the Harmonic Oscillator

Abstract.
A linear quantum harmonic oscillator factors into one dimensional oscillators and can be solved using creation and annihilation operators. We consider a spherical analogue. This analogue does not factor. The two dimensional case is critical, and we compute the spectrum and partition function. This is of interest because the spherical oscillator is potentially relevant to chiral models in 2d quantum field theory.
0. Introduction
A d-dimensional quantum harmonic oscillator is characterized as the unique $SO(d)$ invariant self-adjoint second order differential operator on $\mathbb{R}^{d}$ having principal symbol the Euclidean metric, a Gaussian ground state, and (in a quantum field theoretic context) vanishing ground state energy. The oscillator factors as a sum of independent one dimensional oscillators, and it can be solved using annihilation and creation operators; this is true also of q-deformations of the harmonic oscillator (see e.g. [2]).
Consider the stereographic correspondence
$S^{d}\leftrightarrow\mathbb{R}^{d}\cup\{\infty\}:(x_{1},...,x_{d},y)\leftrightarrow x:=\frac{(x_{1},...,x_{d})}{1-y}$
In the special case $d=1$ with $z=x_{1}+iy\in S^{1}$, this is the Cayley transform
$z=i\frac{x-i}{x+i}\text{ and }x=-i\frac{z+i}{z-i}$
The round metric in stereographic coordinates is given by the formula
$(ds)_{S^{d}}=\frac{2}{1+|x|^{2}}(ds)_{\mathbb{R}^{d}}$
the volume form is given by
$dV_{S^{d}}=2^{d}(1+|x|^{2})^{-d}dV_{\mathbb{R}^{d}}=2^{d}\frac{r^{d-1}}{(1+r^{2})^{d}}drdV_{S^{d-1}}$
and the nonnegative Laplace operator is given by
$\Delta_{S^{d}}=-\frac{1}{4}(1+r^{2})^{2}\left((\frac{d}{dr})^{2}+\frac{1}{r}(1+(d-2)\frac{1-r^{2}}{1+r^{2}})\frac{d}{dr}+\frac{1}{r^{2}}\Delta_{S^{d-1}}\right)$
where $x\in\mathbb{R}^{d}$ and $r=|x|$.
Given $\omega\geq 0$, our spherical analogue of the harmonic oscillator is the differential operator
$L_{\omega}:=\Delta_{S^{d}}+\omega^{2}r^{2}$
This is essentially self-adjoint on compactly supported smooth functions on $\mathbb{R}^{d}$. It is routine to find the spectrum of this operator, it does not factor, and one cannot use annihilation and creation operators. This begs the question of why it is of special interest.
A Gaussian is the ground state of a linear harmonic oscillator. Consider the Cauchy type density $(1+r^{2})^{-\omega}$, relative to $dV_{S^{d}}$, in stereographic coordinates. Is this density the ground state of a natural geometric operator? If $d=2$, then the answer is yes, there is a unique $SO(d)$ invariant self-adjoint second order differential operator having principal symbol determined by the rotationally invariant metric for the d-sphere and having the ground state $(1+r^{2})^{-\omega}$, with vanishing ground state energy:
$\mathbf{L}_{\omega}:=L_{\omega}-\omega=\Delta_{S^{d}}+\omega^{2}r^{2}-\omega$
The story is essentially the same, although slightly more complicated in an interesting way, when we consider the dimension $d$ as a parameter. In general the quadratic potential is quite singular from the point of view of the original $d$-sphere, and breaks the $O(d)$ symmetry down to O(d1)O(d-1). The domain of the operator depends sensitively on $\omega$, e.g. $1$ is in the domain when $\omega=0$, and not when $\omega>0$. It turns out that for the radial eigenfunction equation $r=0$ is regular when $d=1$ and singular otherwise. This is one way to anticipate that various quantities depend discontinuously on the dimension parameter $d$, and $d=2$ is (predictably) critical.
When $d=2$, this Cauchy type ground state arises naturally in connection with the two dimensional quantum sigma model with target a compact semisimple Lie group, i.e. the principal chiral model. It is not known how to rigorously formulate this model, although there is a conjecture for the scattering theory (see [8]). It is possible that the Hamiltonian in finite volume, i.e. when space is a circle $S^{1}$, factors as a sum of ordinary harmonic oscillators and spherical oscillators. We hope to use the partition functions associated to spherical oscillators as a means to test this possibility.
0.1. Plan of the Paper
In Section 1 we recall standard arguments which show that $L_{\omega}$ (with appropriate domain) is a nonnegative self-adjoint operator with compact resolvent (hence discrete spectrum).
In Section 2 we find the ground state and observe how various formulas depend discontinuously on the dimension parameter; $d=2$ is critical.
In Sections 1 and 4 we find the spectrum. We have separated out the $d=2$ case because of its importance and relative simplicity.
In Section 5 we discuss the partition function associated to the operator $\mathbf{L}_{\omega}$. To motivate the paper, we also include a brief outline of the potential relevance of this to the two dimensional principal chiral model.
1. Self-Adjointness
Recall the following three examples from page 331 of [10] of unbounded operators $(T_{k},D(T_{k}))$ in the real Hilbert space $H:=L^{2}([0,1],dx;\mathbb{R})$: $D(T_{1})$ consists of absolutely continuous functions $f$ on $[0,1]$ such that $f^{\prime}\in L^{2}$,
$D(T_{2}):=\{f\in D(T_{1}):f(0)=f(1)\},\text{ and }D(T_{3}):=\{f\in D(T_{1}):f(0)=f(1)=0\}$
and in each case $T_{k}f=\frac{1}{i}f^{\prime}$. Then
$T_{1}^{*}=T_{3},\qquad T_{2}^{*}=T_{2},\qquad T_{3}^{*}=T_{1}$
Heuristically the spectrum of $T_{1}$ consists of all $\lambda\in\mathbb{C}$ with corresponding eigenfunction $f(x):=e^{i\lambda x}$, the spectrum of the self-adjoint operator $T_{2}$ is $2\pi\mathbb{Z}$, and the spectrum of $T_{3}$ is empty. It is important that our spherical oscillators are self-adjoint, to avoid the ”pathology” that occurs in the first and third examples.
An efficient way to define a non-negative self-adjoint operator is to use symmetric forms, see page 11 of [3]. The Hilbert space is $H:=L^{2}(dV_{S^{d}},\mathbb{R})$. The form is
$Q_{\omega}(f,g):=\int(df\wedge*dg+\eta^{2}r^{2}fgdV_{S^{d}})$
where * denotes the star operator for $S^{d}$ (which is conformally invariant precisely when $d=2$), and $Q=Q_{\omega}$ is defined for $f,g$ in the dense subspace $E\subset H$ for which $(Qf,f)$ and $Q(g,g)$ are finite.
Definition 1.
(following [3])
$D(L_{\omega}):=\{g\in E:f\to Q(f,g)\text{ is continuous on }H\}$
and $Q(f,g)=\langle f,L_{\omega}g\rangle_{H}$.
Theorem 1.1.
$L_{\omega}$ is non-negative self-adjoint and the resolvent is compact (more precisely in the Dixmier Schatten class $\mathcal{L}_{d-1}^{+}$).
The first statement follows from the Form Representation Theorem 2.1, page 12 of [3]. Note that $L_{\omega}$ as a differential operator is elliptic in $\mathbb{R}^{d}$ and equal to the self-adjoint $L_{\omega}$ on sufficiently differentiable functions. The second statement follows from the fact that the inverses are pseudo-differential operators of order -2. Note that in the case $d=2$ the inverses are actually in the Dixmier Schatten class $\mathcal{L}_{1}^{+}$.
2. The Ground State
The following shows how the ground state energy depends discontinuously on $\omega$ and the dimension parameter.
Theorem 2.1.
If $\omega>0$, then for any dimension $d$, the ground state for $L_{\omega}$ is
$(1+r^{2})^{-\frac{((d-2)^{2}+16\omega^{2})^{1/2}-(d-2)}{4}}$
and the ground state energy is
$d\frac{((d-2)^{2}+16\omega^{2})^{1/2}-(d-2)}{8}$
where we choose the positive square root in both formulas.
If $\omega=0$, then the above formulas are valid, provided that for $d\leq 2$, we choose the negative square roots, and for $d\geq 2$, we choose the positive square roots.
Proof.
Define $\delta(r):=(1+r^{2})^{-\eta}$, $\eta>0$. Then
$L_{\omega}\delta=\frac{(-2\eta^{2}+(d-2)\eta-2\omega^{2})r^{2}+d\eta}{2}\delta$
For $\delta$ to be an eigenfunction (with eigenvalue $d\eta/2$) it is necessary that $2\eta^{2}+(d-2)\eta-2\omega^{2}=0$, i.e.
$\eta=\frac{-(d-2)\pm((d-2)^{2}+16\omega^{2})^{1/2}}{4}$
This must be nonnegative, vanish when $\omega=0$, and $\delta$ must also be in the domain, in particular
$\int_{0}^{\infty}\delta(r)^{2}(1+r^{2})^{-d}r^{d-1}dr<\infty$
These conditions determine the signs of the square roots that appear in the statement of the theorem. ∎
Remark.
If $d=2$, let $w=1/\overline{z}$, and in general let $w$ denote the reflection of $x$ through the d1d-1 sphere, which we view as a coordinate for $\infty$. In this coordinate the ground state equals
$\frac{|w|^{2\eta}}{(1+|w|^{2})^{\eta}}$
where $\eta=\frac{((d-2)^{2}+16\omega^{2})^{1/2}-(d-2)}{4}$. This vanishes at $\infty$ in the sphere. It is smooth when $\eta$ is integral. Otherwise it has Lipschitz type behavior.
As motivation (see chapter 2 of [1] for the numerous definitions) recall that a Laplace type operator is determined by its principal symbol (the highest order terms), a connection (essentially the first order terms), and a potential. In our narrow context, the latter two conditions can be replaced with invariance, self-adjointness and knowledge of the ground state.
Theorem 2.2.
Suppose $\omega>0$. There is a unique $SO(d)$ invariant self-adjoint second order differential operator having principal symbol determined by the standard metric for the d-sphere and having the ground state $(1+r^{2})^{-\omega}$, with vanishing ground state energy,
$\mathbf{L}_{\omega}=-\frac{1}{4}(1+r^{2})^{2}\left((\frac{d}{dr})^{2}+\frac{1}{r}(1+(d-2)\frac{1-r^{2}}{1+r^{2}})\frac{d}{dr}+\frac{1}{r^{2}}\Delta_{S^{d-1}}\right)$
$d\frac{((d-2)^{2}+16\omega^{2})^{1/2}-(d-2)}{8}$
Proof.
We are led to this formula by the calculation in the preceding proof (with $\eta=\omega$).
Imagine adding lower order terms to $L_{\omega}$. The corresponding quadratic form, in terms of the radial parameter $r$ (we can ignore $\theta$ because of rotational invariance), would (because of self-adjointness) be of the form
$\int_{r=0}^{\infty}b(r)(f^{\prime}(r)g(r)+f(r)g^{\prime}(r))\frac{r}{(1+r^{2})^{2}}dr$
$\int_{r=0}^{\infty}(b(r)f^{\prime}(r)g(r)\frac{r}{(1+r^{2})^{2}}-\frac{\partial}{\partial r}(b(r)f(r)\frac{r}{(1+r^{2})^{2}})dr$
This calculation shows that for the differential operator, the derivative terms cancel, and this only adds zeroth order terms. The fact that we know the ground state implies that $L$ is determined up to a constant. The fact the ground state energy vanishes determines the constant. ∎
3. Spectrum: The Two Dimensional Case
In this section $d=2$.
Theorem 3.1.
Suppose $\omega\geq 0$.
(a) The ground state for $L_{\omega}$ is
$(1+r^{2})^{-\omega}$
with corresponding eigenvalue $\omega$.
(b) The spectrum is
$\lambda_{m,n}=(m+\frac{1}{2}(n+1+\sqrt{n^{2}+4\omega^{2}}))^{2}-\omega^{2}-\frac{1}{4}$
$m,n\geq 0$, with multiplicity $\mathbf{m}(m,n)=1$ when $n=0$ $\mathbf{m}(m,n)=2$ otherwise (The values of the $\lambda_{m,n}$ are not necessarily distinct, see part (c) and the examples in Remark Remark below). The corresponding eigenfunctions are of the form
$\frac{r^{n}}{(1+r^{2})^{(\sqrt{4\omega^{2}+4\lambda+1}-1)/2}}F(-m-\sqrt{4\omega^{2}+n^{2}},-m,n+1,-r^{2})f(\theta)$
where $F$ is the hypergeometric function (an even $2m$ degree polynomial in $r$ in our case) and $f(\theta)$ is a linear combination of $cos(n\theta)$ and $sin(n\theta)$.
(c) If $\omega=0$, then the spectrum is $N(N+1)$ with multiplicity $2N+1$, $N=0,1,$....
Remark.
In reference to part (b), note that if $\omega=2\sqrt{3}$, then $\lambda_{3,1}=44=\lambda_{1,4}$, and if $\omega=4$, then $\lambda_{9,6}=\lambda_{1,15}$. Thus degeneracy of the eigenvalues can definitely occur, even if $\omega\neq 0$. Nonetheless in part (b) we are counting the eigenvalues correctly. When $\omega=2\sqrt{3}$, then $\lambda_{3,1}$ and $\lambda_{1,4}$ are the same, but if we attach a multiplicity $2$ to each of them, it is the same as attaching multiplicity $4$ to the common value. This trivial point is important when we consider the partition function in Section 5 below.
We have not been able to find a robust deterministic condition which rules out degeneracy.
Proof.
We have already proven (a) (We will add an abstract geometric argument in Remark Remarks following the proof, to explain how we were led to the operator from the ground state).
To find eigenvalues and eigenfunctions, we use separation of variables (we will recall the standard justification at the end of the proof). Suppose we consider an eigenfunction of the form $g(r)f(\theta)$. Then
$\frac{1}{4}(1+r^{2})^{2}\left(g^{\prime\prime}(r)f(\theta)+\frac{1}{r}g^{\prime}(r)f(\theta)+\frac{1}{r^{2}}g(r)f^{\prime\prime}(\theta)\right)+\omega^{2}r^{2}g(r)f(\theta)=\lambda g(r)f(\theta)$
Then
$(1+r^{2})^{2}\left(\frac{g^{\prime\prime}(r)}{g(r)}+\frac{1}{r}\frac{g^{\prime}(r)}{g(r)}+\frac{1}{r^{2}}\frac{f^{\prime\prime}(\theta)}{f(\theta)}\right)+4\omega^{2}r^{2}=4\lambda$
and
$\frac{r^{2}}{(1+r^{2})^{2}}\left(-(1+r^{2})^{2}(\frac{g^{\prime\prime}(r)}{g(r)}+\frac{1}{r}\frac{g^{\prime}(r)}{g(r)})+4\omega^{2}r^{2}-4\lambda\right)=\frac{f^{\prime\prime}(\theta)}{f(\theta)}$
This implies that $f(\theta)$ is a combination of $cos(n\theta)$ and $sin(n\theta)$, where $n$ is a nonnegative integer, and
(3.1) $r^{2}g^{\prime\prime}(r)+rg^{\prime}(r)+\left(-n^{2}+4(1+r^{2})^{-2}(-\omega^{2}r^{4}+\lambda r^{2})\right)g(r)=0$
This has regular singular points and is therefore equivalent to the hypergeometric equation. The indicial equation is $\alpha^{2}-n^{2}=0$ and $g$ has the form
$g(r)=r^{\pm n}\sum_{l=0}^{\infty}g_{l}r^{l}$
The relevant solution involves +n+n, so that $g$ is regular at $r=0$. One finds (using Maple)
(3.2) $g(r)=\frac{r^{n}}{(1+r^{2})^{(\sqrt{4\omega^{2}+4\lambda+1}-1)/2}}F(a,b,n+1,-r^{2})$
where $F$ denotes the hypergeometric function with
$a=\frac{1}{2}(-\sqrt{4\omega^{2}+4\lambda+1}+n+1-\sqrt{4\omega^{2}+n^{2}})$
and
$b=\frac{1}{2}(-\sqrt{4\omega^{2}+4\lambda+1}+n+1+\sqrt{4\omega^{2}+n^{2}})=a+\sqrt{4\omega^{2}+n^{2}}$
A necessary condition for $\lambda$ to be an eigenvalue is that $g(r)$ is square integrable on $(0,\infty)$ with respect to $(1+r^{2})^{-2}rdr$. Thus $a$ or $b$ must be a non-positive integer. As above $a<b$.
Examples.
Suppose $\omega=0$. In this case $b=a+n$. Suppose b=mb=-m, a non-positive integer. In this case $\lambda=\lambda_{m,n}=(m+n)(m+n+1)$. Square integrability holds. Now consider the possibility that a=ma=-m, a non-positive integer and b=a+n=m+n>0b=a+n=-m+n>0, i.e. $n>m$. In this case $\lambda=m(m+1)$. For $g(r)$ to be square integrable it is necessary that
$2(\sqrt{1+4\lambda}-1)-2n-4m+3>1$
or
$1+4\lambda=1+4m(m+1)>(n+2m)^{2}=n^{2}+4mn+4m^{2}$
or
$1+4m(1)>(n+2m)^{2}=n^{2}+4mn$
Square integrability fails. So we need both a and b to be non-positive. If we fix $\lambda=N(N+1)$, then the multiplicity is $2N+1$. For if $\lambda=\lambda_{m,n}$, when $n=0$ we can only choose $cos(n\theta)=1$ and when $n>0$ we can choose $cos(n\theta)$ or $sin(n\theta)$.
Now suppose $\omega>0$.
Suppose that b=mb=-m, a non-positive integer. Since $a<b$, $F(a,b,c,-r^{2})$ is a polynomial of degree $m$ in $z=-r^{2}$. In this case (after a remarkable simplification)
$\lambda:=\lambda_{m,n}=(m+\frac{1}{2}(n+1+\sqrt{n^{2}+4\omega^{2}}))^{2}-\omega^{2}-\frac{1}{4}$
as in the statement of the Theorem. We first find the condition for $g\in H:=L^{2}((1+r^{2})^{-2}rdr)$. Since $F(a,b,n+1,-r^{2})$ is a polynomial of degree $2m$ in $r$, $g\in L^{2}$ if and only if (using the expression (3.2) for the eigenfunction)
$2(\sqrt{4\omega^{2}+4\lambda+1}-1)-2n-4m)+3>1$
or
$4\omega^{2}+4\left((m+\frac{1}{2}(n+1+\sqrt{n^{2}+4\omega^{2}}))^{2}-\omega^{2}-\frac{1}{4}\right)+1>(n+2m)^{2}$
or
$4(m+\frac{1}{2}(n+1+\sqrt{n^{2}+4\omega^{2}}))^{2}>(n+2m)^{2}$
This condition is always satisfied, so $g(r)f(\theta)$ is in the Hilbert space $H$. To see that it is actually in $D(L_{\omega})$, for $F\in C^{\infty}(\mathbb{R}^{2})$ having compact support, we use integration by parts. This is justified, because the functions are smooth in the plane and $F$ vanishes outside of a compact set. Consequently
$\int(dF\wedge*d(gf)+\omega^{2}r^{2}Fgf)=\int F((\Delta+\omega^{w}r^{2})(gf))dV_{S^{2}}=\langle F,\lambda gf\rangle_{H}$
This shows that the function $H\to\mathbb{R}:F\to Q(F,gf)$ is continuous, hence $gf\in D(L_{\omega})$. Thus $\lambda_{m,n}$ is actually an eigenvalue.
Suppose that a=ma=-m and $b=a+\sqrt{4\omega^{2}+n^{2}}>0$. In this case
$\lambda=\lambda_{m,n}^{\prime}=(m+\frac{1}{2}(n+1-\sqrt{n^{2}+4\omega^{2}}))^{2}-\omega^{2}-\frac{1}{4}$
(the sign of the square root has changed). Again $F(a,b,n+1,-r^{2})$ is a polynomial of degree $2m$ in $r$. Thus $g\in L^{2}$ if and only if
$2(\sqrt{4\omega^{2}+4\lambda+1}-1)-2n-4m+3>1$
or
$4\omega^{2}+4\left((m+\frac{1}{2}(n+1-\sqrt{n^{2}+4\omega^{2}}))^{2}-\omega^{2}-\frac{1}{4}\right)+1>(n+2m)^{2}$
or
$4(m+\frac{1}{2}(n+1-\sqrt{n^{2}+4\omega^{2}}))^{2}>(n+2m)^{2}$
This implies $1>\sqrt{n^{2}+4\omega^{2}}$, hence $n=0$. But then $b=-m+2\omega\leq 0$, which contradicts our assumption that $b>0$. Thus $\lambda_{m,n}^{\prime}$ is not an eigenvalue.
We briefly recall why separation of variables yields a complete set of eigenfunctions. Suppose that $\psi$ is an eigenfunction corresponding to an eigenvalue $\lambda$. By ellipticity $\psi$ is smooth on $\mathbb{R}^{2}$ and will have an expansion of the form
$\psi(r,\theta)=\sum_{n=1}^{\infty}(\psi_{n}(r)cos(n\theta)+\widetilde{\psi_{n}}(r)sin(n\theta)$
This implies
$L_{\omega}\psi=\sum_{n=1}^{\infty}\left((-\frac{1}{4}(1+r^{2})^{2}\left(\psi_{n}^{\prime\prime}(r)+\frac{1}{r}psi_{n}^{\prime}(r)-\frac{1}{r^{2}}n^{2}\psi_{n}(r)\right)+\omega^{2}r^{2}\psi_{n}(r)\right)cos(n\theta)$
plus a similar sum involving $\widetilde{\psi}_{n}$ and sine functions. This equals $\lambda\psi$. This implies that for each $n$
$(-\frac{1}{4}(1+r^{2})^{2}\left(\psi_{n}^{\prime\prime}(r)+\frac{1}{r}psi_{n}^{\prime}(r)-\frac{1}{r^{2}}n^{2}\psi_{n}(r)\right)+\omega^{2}r^{2}\psi_{n}(r)=\lambda\psi_{n}(r)$
and similarly for the $\widetilde{\psi}_{n}$. Since we found all of the solutions for the radial eigenfunction problem (3.1), this implies that $\lambda$ is equal to one of the $\lambda_{m,n}$.
This completes the proof.
Remarks.
(a) When $d=2$, the ground state calculation, and the origin of the operator $L_{\omega}$, can be cast in more abstract terms. Suppose that $L$ is a holomorphic Hermitian line bundle with canonical connection $\nabla$, where $\nabla^{(0,1)}=\overline{\partial}$. If $s$ is a holomorphic section, then
$\partial(s,s)=\theta(s,s)$
and
$\overline{\partial}\partial(s,s)=\overline{\partial}(\theta)(s,s)-\theta\wedge\overline{\theta}(s,s)$
(see page 73 of [6]). Thus
$\overline{\partial}\partial(s,s)+\theta\wedge\overline{\theta}(s,s)=\Theta(s,s)$
where $\Theta$ is the curvature, a (1,1) form. In our case the line bundle is the kth power of the dual of the canonical bundle (or the 2kth power of the generating positive line bundle on $\mathbb{P}^{1}$, the dual of the tautological bundle). The section $s$ is a canonical (highest weight) holomorphic section such that $(s,s)=(1+r^{2})^{-2k}$ and the curvature is a multiple (essentially the ground state energy) of the area form on the base (because of homogeneity).
(b) This point of view leads to one possible generalization, involving Kahler-Einstein manifolds (where the Ricci form is a multiple of the symplectic form on the base).
A related observation is that stereographic coordinates is a special case of Cayley transform coordinates for classical symmetric spaces. This leads to other generalizations.
4. Spectrum: The General Case
If $\omega=0$, then it is well-known that the spectrum of the positive Laplacian on the $d$-sphere is N(N+d1)N(N+d-1) with multiplicity
$\left(\begin{matrix}(N+d\\ N\end{matrix}\right)-\left(\begin{matrix}(N+d-2\\ d\end{matrix}\right),\qquad N=0,1,$...
In particular the ground state energy is zero.
Theorem 4.1.
Suppose that $\omega>0$. For $L_{\omega}$
(a) The ground state is
$\delta(r):=(1+r^{2})^{-\frac{|((d-2)^{2}+16\omega^{2})^{1/2}|-(d-2)}{4}}$
with ground state energy
$d\frac{|((d-2)^{2}+16\omega^{2})^{1/2}|-(d-2)}{8}$
(b) If $d\geq 2$, then the spectrum is
$\lambda_{m,n}=(m+\frac{1}{4}(|((2n+d-2)^{2}+16\omega^{2})^{1/2}|+(d-2+2(n+1)))^{2}-\omega^{2}-\frac{1}{4}(d-1)^{2}$
$m,n\geq 0$, with multiplicity $\mathbf{m}(m,n)=1$ when $n=0$ and
$\mathbf{m}(m,n)=\left(\begin{matrix}N+d-1\\ N\end{matrix}\right)-\left(\begin{matrix}N+d-3\\ N-3\end{matrix}\right)$
when $n>0$, where N=n(n+d2)N=n(n+d-2) (As in the $d=2$ case, these values are not necessarily distinct, which happens for example when $\omega=0$).
(b)’ If $d=1$, then the spectrum is $\lambda_{m}:=\lambda_{m,n=0}$, $m\geq 0$, where we choose the negative square root in the formula for $\lambda_{m,n=0}$:
$\lambda_{m}=(m+\frac{1}{4}(1-|(1+16\omega^{2})^{1/2}|))^{2}-\omega^{2},\qquad m=0,1,$...
If $\omega>0$, the spectrum is multiplicity free.
Proof.
We proved (a) in Section 2 and observed this depends discontinuously on $\omega$ and the dimension.
(b) Assume $d>1$. To find eigenvalues and eigenfunctions, we use separation of variables. Suppose we consider an eigenfunction of the form $g(r)f(\theta)$, where (abusing notation) we are writing $f(\theta)$ for a function that is defined on the d1d-1 sphere. Then
$\frac{1}{4}(1+r^{2})^{2}\left(g^{\prime\prime}(r)f(\theta)+\frac{1}{r}(1+(d-2)\frac{1-r^{2}}{1+r^{2}})g^{\prime}(r)f(\theta)+\frac{1}{r^{2}}g(r)\Delta_{S^{d-1}}f(\theta)\right)$
$\omega^{2}r^{2}g(r)f(\theta)=\lambda g(r)f(\theta)$
Thus
$(1+r^{2})^{2}\left(\frac{g^{\prime\prime}(r)}{g(r)}+\frac{1}{r}(1+(d-2)\frac{1-r^{2}}{1+r^{2}})\frac{g^{\prime}(r)}{g(r)}+\frac{1}{r^{2}}\frac{\Delta_{S^{d-1}}f(\theta)}{f(\theta)}\right)+4\omega^{2}r^{2}=4\lambda$
and
$\frac{r^{2}}{(1+r^{2})^{2}}\left(-(1+r^{2})^{2}(\frac{g^{\prime\prime}(r)}{g(r)}+\frac{1}{r}(1+(d-2)\frac{1-r^{2}}{1+r^{2}})\frac{g^{\prime}(r)}{g(r)})+4\omega^{2}r^{2}-4\lambda\right)=\frac{\Delta_{S^{d-1}}f(\theta)}{f(\theta)}$
This implies that $f(\theta)$ is an eigenfunction for $\Delta_{S^{d-1}}$. The corresponding eigenvalue is of the form n(n+d2)n(n+d-2), where $n$ is a nonnegative integer, and
$r^{2}g^{\prime\prime}(r)+(1+(d-2)\frac{1-r^{2}}{1+r^{2}})rg^{\prime}(r)+\left(-n(n+d-2)+4(1+r^{2})^{-2}(-\omega^{2}r^{4}+\lambda r^{2})\right)g(r)=0$
This has regular singular points and is equivalent to the hypergeometric equation. At $r=0$ the indicial equation is
$\alpha^{2}+(d-2)\alpha-n(n+d-2)=0$
for which the roots are $\alpha=n$ and $\alpha=-(n+d-2)$. For $g(r)$ to be regular at $r=0$, $g(r)$ must be of the form
$g(r)=r^{n}\sum_{l=0}^{\infty}g_{l}r^{l}$
Using Maple,
(4.1) $g(r)=r^{n}\left(1+r^{2}\right)^{\frac{d-1-\sqrt{(d-1)^{2}+4\omega^{2}+4\lambda}}{2}}F(a,b,n+\frac{d}{2},-r^{2})$
where $F$ denotes the hypergeometric function, we choose the positive square root, and
$a=-\frac{\sqrt{(d-1)^{2}+4\omega^{2}+4\lambda}}{2}+\frac{n}{2}+\frac{d}{4}-\frac{\sqrt{(2n+d-2)^{2}+16\omega^{2}}}{4}$
$b=-\frac{\sqrt{(d-1)^{2}+4\omega^{2}+4\lambda}}{2}+\frac{n}{2}+\frac{d}{4}+\frac{\sqrt{(2n+d-2)^{2}+16\omega^{2}}}{4}$
For $\lambda$ to be an eigenvalue, $g(r)$ must be square integrable on $(0,\infty)$ with respect to $(1+r^{2})^{-d}r^{d-1}dr$. A necessary condition is that $a$ or $b$ is a non-positive integer.
Suppose that a<b=ma<b=-m, a non-positive integer. In this case $F(a,b,n+\frac{d}{2},-r^{2})$ is a polynomial of degree $m$ in $r^{2}$. In this case $\lambda=\lambda_{m,n}$ equals (after a somewhat miraculous simplification)
$(m+\frac{1}{4}((2n+d-2)^{2}+16\omega^{2})^{1/2}+(d-2+2(n+1)))^{2}-\omega^{2}-\frac{1}{4}(d-1)^{2}$
For $\lambda_{m,n}$ to be an eigenvalue it is necessary that $g\in L^{2}((1+r^{2})^{-d}r^{d-1}dr)$. Now $F(a,b,n+\frac{d}{2},-r^{2})$ is a polynomial of degree $2m$ in $r$. Thus $g\in L^{2}$ if and only if
$2n+2(d-1-\sqrt{(d-1)^{2}+4\omega^{2}+4\lambda_{m,n}})+4m+d-1-2d<-1$
or
$2n+4m+d-2<2\sqrt{(d-1)^{2}+4\omega^{2}+4\lambda_{m,n}}$
or
$(2n+4m+d-2)^{2}<4((d-1)^{2}+4\omega^{2}+4\lambda_{m,n})$
or
$(2n+4m+d-2)^{2}<4(m+\frac{1}{4}(((2n+d-2)^{2}+16\omega^{2})^{1/2}+(d-2+2(n+1)))^{2}$
This condition is always satisfied, so $g(r)f(\theta)$ is in the Hilbert space. We use the same integration by parts argument as in the $d=2$ case to see that $g(r)f(\theta)\in D(L_{\omega})$, which implies $\lambda_{m,n}$ is an eigenvalue.
As in the $d=2$ case we rule out $\lambda$ as an eigenvalue when a=ma=-m and $b>0$.
Separation of variables yields a complete set of eigenfunctions by the same argument as in the proof of the $d=2$ case.
Suppose that $d=1$. In this case we are not restricting $r$ to be positive. The eigenvalue equation
$g^{\prime\prime}(r)+\frac{2r}{1+r^{2}}g^{\prime}(r)+4(1+r^{2})^{-2}(-\omega^{2}r^{2}+\lambda)g(r)=0$
no longer has a singularity at $r=0$, and
$g(r)=c_{1}LegendreP(\sqrt{16\omega^{2}+1}/2-1/2,2\sqrt{\omega^{2}+\lambda},ir)$
$c_{2}LegendreQ(\sqrt{16\omega^{2}+1}/2-1/2,2\sqrt{\omega^{2}+\lambda},ir)$
where the signs of the square roots have to be determined.
For each $m=0,$... we find that
$LegendreQ(-|\sqrt{16\omega^{2}+1}|/2-1/2,-2|\sqrt{\omega^{2}+\lambda_{m}}|,ir)$
$C(1+r^{2})^{-|\sqrt{\omega^{2}+\lambda_{m}}|}F(-m-\frac{1}{2},-m,1-|\sqrt{\omega^{2}+\lambda_{m}}|,ir)$
is an eigenfunction corresponding to $\lambda_{m}=(m+\frac{1}{4}(1-|(1+16\omega^{2})^{1/2}|)^{2}-\omega^{2}$. Using
$LegendreP(a,b,z)=C\left(\frac{z+1}{z-1}\right)^{b/2}F(-a,a+1,1-b,1/2-z/2))/((z-1)^{(}b/2)$
which implies
$|LegendreP(\sqrt{16\omega^{2}+1}/2-1/2,2\sqrt{\omega^{2}+\lambda_{m}},ir)|$
$C|F(-m-\frac{1}{2},-m,1-|\sqrt{\omega^{2}+\lambda_{m}}|,ir)$
we can rule out that LegendreP is an eigenfunction, because this is not square integrable, unless $m=0$. Hence the spectrum is multiplicity free. ∎
Corollary 1.
Suppose that $d\geq 1$, $\omega\geq 0$, and let $\mathbf{L}_{\omega}$ denote $L_{\omega}$ minus its ground state energy.
(a) The ground state is
$\delta(r):=(1+r^{2})^{-\frac{d-2+|((d-2)^{2}+16\omega^{2})^{1/2}|}{4}}$
(b) If $d\geq 2$, then the spectrum is
$\lambda_{m,n}=(m+\frac{1}{4}(|((2n+d-2)^{2}+16\omega^{2})^{1/2}|+(d-2+2(n+1)))^{2}-(\frac{1}{4}(|((d-2)^{2}+16\omega^{2})^{1/2}|+d))^{2}$
$m,n\geq 0$, with multiplicity $\mathbf{m}(m,n)=1$ when $n=0$ and
$\mathbf{m}(m,n)=\left(\begin{matrix}(N+d-1\\ N\end{matrix}\right)-\left(\begin{matrix}(N+d-3\\ N-3\end{matrix}\right)$
when $n>0$, where N=n(n+d2)N=n(n+d-2).
(b)’ If $d=1$, then the spectrum is
$\lambda_{m}=(m+\frac{1}{4}(1-|(1+16\omega^{2})^{1/2}|))^{2}-(\frac{1}{4}(1-|(1+16\omega^{2})^{1/2}|))^{2}$
$m\geq 0$, with multiplicity $1$.
5. The Partition Function
In this section we will briefly discuss the partition function, and why it might be relevant to the principal chiral model in two dimensions.
As in Corollary 1, let $\mathbf{L}_{\omega}$ denote $L_{\omega}$ minus its ground state energy. The partition function is $tr(e^{-t\mathbf{L}_{\omega}})$, $t>0$. For example when $d=2$,
(5.1) $tr(e^{-t\mathbf{L}_{\omega}})=\sum_{m,n\geq 0}\mathbf{m}(m,n)e^{-t(\lambda_{m,n}-\omega)}$
$e^{t(\omega+\frac{1}{2})^{2}}\sum_{m,n\geq 0}\mathbf{m}(m,n)e^{-t(m+\frac{n+1+\sqrt{n^{2}+4\omega^{2}}}{2})^{2}}$
where $\mathbf{m}(m,n)=1$ when $n=0$ and =2=2 when $n>0$. For any dimension $d\geq 1$ and $\omega\geq 0$, the partition function has an asymptotic expansion
(5.2) $tr(e^{-t\mathbf{L}_{\omega}})\sim(4\pi t)^{-d/2}vol(S^{d})\sum_{j=0}^{\infty}a_{i}(\omega)t^{i}$
(see e.g. Section 1.10 of [5]). This trace formula is crucial in understanding the zeta function and determinant associated to $\mathbf{L}_{\omega}$, but we will not pursue this.
5.0.1. Example: $d=1$
Suppose that $d=1$. If $\omega=0$, then $\mathbf{L}_{\omega=0}=\Delta_{S^{1}}$ and
$tr(e^{-t\mathbf{L}_{\omega=0}})=\sum_{n=-\infty}^{\infty}e^{-tn^{2}}$
Poisson summation implies this equals
$\pi^{1/2}t^{-1/2}\sum_{n=-\infty}^{\infty}e^{-\pi^{2}n^{2}/t}=\pi^{1/2}t^{-1/2}(1+2\sum_{n=1}^{\infty}e^{-\pi^{2}n^{2}/t})$
The exponential smallness of the terms in the sum implies the asymptotic expansion
$tr(e^{-t\mathbf{L}_{\omega=0}})\sim\pi^{1/2}t^{-1/2}$
to any order in powers of $t>0$, meaning that for any $N>0$ there is a constant $c_{N}$ such that
$t^{1/2}tr(e^{-tL_{\omega=0}})<c_{N}t^{N}\text{ for }0<t<1$
If $\omega>0$, then (b)’ of Corollary 1 implies
$tr(e^{-t\mathbf{L}_{\omega}})=\sum_{m=0}^{\infty}e^{-t((m+\frac{1}{4}(1-|\sqrt{1+16\omega^{2}}|))^{2}-(\frac{1}{4}(1-\sqrt{|1+16\omega^{2}|}))^{2}}$
This family of functions can be parameterized by $m_{0}=m_{0}(\omega):=\frac{1}{4}(|\sqrt{1+16\omega^{2}}|-1)\geq 0$. In terms of this parameter
(5.3) $tr(e^{-t\mathbf{L}_{\omega}})=e^{tm_{0}^{2}}\sum_{m=0}^{\infty}e^{-t(m-m_{0})^{2}}$
These are shifts of partial theta functions; there is probably not much more we can say. Since this is dominated term by term by the partition function for $\omega=0$, this has an asymptotic expansion in powers of $t>0$ which is independent of $\omega$.
5.0.2. Example: $d=2$
If $\omega=0$, then
(5.4) $tr(e^{-tL_{\omega=0}})=\sum_{n=0}^{\infty}(2n+1)e^{-t(n^{2}+n)}=e^{t/4}\sum_{n=0}^{\infty}(2n+1)e^{-t(n+\frac{1}{2})^{2}}$
This resembles the derivative of a theta function, but there does not seem to be an exact formula in terms of theta functions.
It is remarkable that Mulholland ([7]) found an exact asymptotic expansion for the sum in (5.5).
Proposition 1.
There is an asymptotic expansion
$\sum_{n=0}^{\infty}(2n+1)e^{-t(n+\frac{1}{2})^{2}}\sim\frac{1}{t}+a_{0}(0)+a_{1}(0)t+$...
where
$a_{n}(0)=\frac{(-1)^{n}}{n+1}B_{2n+2}(1-2^{-2n-1})$
and the $B_{m}$ are the Bernoulli numbers:
$B_{2n+2}=(-1)^{n}\frac{2(2n+2)!\zeta(2n+2)}{(2\pi)^{2n+2}}$
This asymptotic series is not convergent, hence it does not determine the original sum (Mulholland observes that the sum of odd (or even) terms has the same asymptotic series). It is striking that the $a_{n}(0)$ are positive.
When $\omega>0$ the partition function
(5.5) $tr(e^{-t\mathbf{L}_{\omega}})=\sum_{m,n\geq 0}\mathbf{m}(m,n)e^{-t(\lambda_{m,n}-\omega)}$
$e^{t(\omega+\frac{1}{2})^{2}}\sum_{m,n\geq 0}\mathbf{m}(m,n)e^{-t(m+\frac{n+1+\sqrt{n^{2}+4\omega^{2}}}{2})^{2}}$
is considerably more complicated, because it is a double sum. Unlike the case $d=1$ there does not seem any hope of expressing it in terms of standard functions. Numerical experiments seem to show that in the asymptotic expansion (5.2) the $a_{n}(\omega)$ are increasing functions of $\omega$. But we do not see how to improve on general formulas.
5.1. Relation to the Principal Chiral Model
In this subsection we will very briefly outline why it is possible that the Hamiltonian for the quantum principal chiral model might be a sum of ordinary harmonic oscillators and spherical harmonic oscillators; see Section 5 of [9] for more details.
If $X$ is a Riemannian manifold, then the action for the sigma model with target $X$ on the cylinder $S^{1}\times\mathbb{R}$ is
(5.6) $S(x:S^{1}\times\mathbb{R}\to X)=\frac{1}{2}\int_{\Sigma}\left(|\frac{\partial x}{\partial t}|^{2}+|\frac{\partial x}{\partial\theta}|^{2}\right)d\theta dt$
The time zero fields constitute the loop space $Map(S^{1},X)$.
If $X=\mathbb{R}$, this is a massless free scalar field. In terms of the Fourier series expansion $x(t,\theta)=\sum_{n=-\infty}^{\infty}x_{n}(t)e^{in\theta}$, the action is
$S(x)=\sum_{n=-\infty}^{\infty}\int_{t}(|x_{n}^{\prime}(t)|^{2}+n^{2}|x_{n}(t)|^{2})$
which is quadratic. Consequently the Feynman measure for the free field is Gaussian. The ground state of the theory is obtained by taking the trace on $S^{1}$. This is also a Gaussian measure, and it factors in terms of the Fourier modes $x_{n}$, $n=1,2,$.... From this we can deduce the form of the Hamiltonian for the free field: it is a sum of ordinary harmonic oscillators (all of this is well-known; see pages 733-734 of [4] for a lucid discussion).
Now suppose that $X$ is a simply connected compact Lie group with the bi-invariant metric; we assume $X=SU(2)$ (the 3-sphere) for definiteness. In this case the action does not factor in a simple way, and it is necessary to use renormalization group techniques to even understand the model at an effective field theory level (see Lecture 3 of [4]). We will persist from a geometric point of view.
The first problem is to understand the state space for the sigma model. At a heuristic level, this should be an $L^{2}$ space over the space of time zero fields, i.e. the loop group $Map(S^{1},SU(2,\mathbb{C}))$. It is remarkable that there exists a kind of Haar measure (denoted by $\mu_{0}$) which lives on a completion of the loop group (the hyperfunction completion of the loop group, see [9]). There are reasons to believe this is the ground state for the quantum model. There is also a noncommutative, or multiplicative, analogue of Fourier series: Given a $g\in W^{1/2,L^{2}}(S^{1},K)$ (the critical degree of smoothness for a loop, in the $L^{2}$ Sobolev sense), there is a unique ‘multiplicative Fourier series’ (with $z=exp(i\theta)$
(5.7) $g(z)=\prod_{i\geq 0}^{\rightarrow}\mathbf{a}(\eta_{i})\left(\begin{matrix}1&\overline{\eta}_{i}z^{i}\\ -\eta_{i}z^{-i}&1\end{matrix}\right)\left(\begin{matrix}e^{\chi(z)}&0\\ 0&e^{-\chi(z)}\end{matrix}\right)\prod_{k\geq 1}^{\leftarrow}\mathbf{a}(\zeta_{k})\left(\begin{matrix}1&\zeta_{k}z^{-k}\\ -\overline{\zeta}_{k}z^{k}&1\end{matrix}\right)$
where $\chi(z)=\sum\mathbf{\chi}_{j}z^{j}$ is a $i\mathbb{R}$-valued Fourier series (modulo $2\pi i\mathbb{Z}$). Here the $\eta_{i}$ and $\zeta_{j}$ variables should be viewed as affine coordinates for 2-spheres.
Although $W^{1/2,L^{2}}(S^{1},K)$ has measure zero with respect to $\mu_{0}$, the $\eta_{i}$, $\chi_{j}$ and $\zeta_{k}$ can be interpreted as random variables with respect to $\mu_{0}$, and conjecturally $\mu_{0}$ factors:
$d\mu_{0}=\left(\prod_{i=0}^{\infty}\frac{1+2i}{\pi}\frac{d\lambda(\eta_{i})}{(1+|\eta_{i}|^{2})^{2+2i}}\right)\times\left(\prod_{j=1}^{\infty}\frac{4j}{\pi}e^{-4j|\mathbf{\chi}_{j}|^{2}}d\lambda(\mathbf{\chi}_{j})\right)$
(5.8) $\times d\lambda(e^{\chi_{0}})\times\left(\prod_{k=1}^{\infty}\frac{2k-1}{\pi}\frac{d\lambda(\zeta_{k})}{(1+|\zeta_{k}|^{2})^{2k}}\right),$
This form of the vacuum, involving a product of Cauchy type densities and Gaussians, suggests a hypothesis for the form for the Hamiltonian. The Gaussian density for $\chi$ (and asymptotic freedom) suggests that $\chi$ is a massless free field. The partition function can be expressed in terms of the $\eta$ function:
$trace(e^{-tH_{\chi}})=\frac{1}{\eta(4\tau)^{2}}\text{ where }\tau=\frac{it}{2\pi}$
The other variables appear to involve spherical harmonic oscillators, and we have computed the partition functions for these. If these oscillators are independent, which is a wild possibility, then the full partition function is
$trace(e^{-tH(1)})=\frac{\left(e^{t}\prod_{k=1}^{\infty}\sum_{m,n=0}\mathbf{m}(m,n)e^{-t(\lambda_{k,m,n}-2k)}\right)^{2}}{\eta(e^{-4t})^{2}}$
where
$\lambda_{k,m,n}=m^{2}+m+m(n+\sqrt{n^{2}+16k^{2}})+(n+1)\frac{n+\sqrt{n^{2}+16k^{2}}}{2}-2k$
and $\mathbf{m}(m,n)=1$ when $n=0$ and =2=2 otherwise.
We have written this speculative formula for the Hamiltonian as $H(1)$ because we are assuming the radius of the circle is one. The Hamiltonian should depend on the radius, which we do not fully understand (hence we do not know how to compare the spectrum of $H(1)$ with the conjectural scattering solution in [8]). When the radius is small the conjectured asymptotic freedom of the model should mean the theory behaves like a massless free field, i.e. that the $\chi$ degrees of freedom, which are Gaussian, should dominate. When the radius is large, i.e. when we are considering the model in the plane, then the other variables should dominate, and we should be able to use the spherical oscillators to reproduce other expected properties of the model.
In any event we now have a possible formula for the partition function, and we are seeking a numerical test for it.